OpenStudy (anonymous):

What are the roots of the quadratic equation xÆ – 2x + 4 = 8x – 21?

OpenStudy (jdoe0001):

do you know how to factor a quadratic?

OpenStudy (anonymous):

No

OpenStudy (jdoe0001):

ok... so.. .what instructions are you meant to use to get the factors?

OpenStudy (anonymous):

im sorry i dont understand ..

OpenStudy (jdoe0001):

well.. you're expected to find the roots, or zeros what instructions are you meant to use to find the roots ? there are several ways... but dunno what the exercise is covering

OpenStudy (anonymous):

I mean i honestly dont know. Im taking an online math class and its not really teaching me anything. It just gives me problems like these and thats it .

OpenStudy (jdoe0001):

hmmm well... sounds like they're not help then you're meant to have covered that section before doing the exercises otherwise..... I mean.. .what if decide to give you an exercise on say physics in initial velocity, if you dunno what that is.or haven't covered it.. the exercise wouldn't apply

OpenStudy (anonymous):

Yea there really not. And its awful because i need this class to graduate. Its awful and its hard. Thats why im on this website to get help

OpenStudy (jdoe0001):

well.... do you know how to do linear simplification? like say, solving for a variable

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

ok ... then let's see yours \(\bf x^2 - 2x + 4 = 8x - 21\) if you moved everything to the left-hand-side.... what would you be left with?

OpenStudy (jdoe0001):

bearing in mind that the "x" and constant terms, simplify themselves

OpenStudy (anonymous):

x^2-2x-8x+4-21=0 .. ?

OpenStudy (anonymous):

Is there anyway you can help me with my other assignments ?

OpenStudy (jdoe0001):

well \(\large { x^2 - 2x + 4 = 8x - 21\implies x^2-2x+4-8x+21=0 \\ \quad \\ x^2-10x+25=0\qquad notice\implies \begin{array}{cccllll} x^2&-10x&+25&=0\\ &\uparrow &\uparrow \\ &-5-5&-5\cdot -5 \end{array} }\) notice that, the middle term can come from ADDING 2 factors, of the last term "25" and that the same factors can be used to produce the last term, or "25"